OpenStudy (maarriinnaaaa):

Evaluate 32 + (7 − 2) ⋅ 4 − 6 over 3

rebeccaxhawaii (rebeccaxhawaii):

MM

OpenStudy (maarriinnaaaa):

@rebeccaxhawaii ???????????

OpenStudy (anonymous):

\[\frac{1}{3} (((7-2) 4-6)+32)=\frac{46}{3} \]

OpenStudy (anonymous):

\[\frac{ 32 + 5⋅4 - 6 }{ 3 }\] = \[\frac{ 32 + 20 - 6 }{ 3 }\] = \[\frac{ 46 }{ 3 }\] = \[15.33333333333333\]